arthham

JEE Advanced 2024 · Probability (Class 12)

A student appears for a quiz consisting of only true-false type questions and answers all the questions. The student knows the answers of some questions and guesses the answers for the remaining questions. Whenever the student knows the answer of a question, he gives the correct answer. Assume that the probability of the student giving the correct answer for a question, given that he has guessed it, is 12\frac{1}{2}. Also assume that the probability of the answer for a question being guessed, given that the student's answer is correct, is 16\frac{1}{6}. Then the probability that the student knows the answer of a randomly chosen question is

Answer: C. 57\frac{5}{7}

Watch this solved on the board The full walkthrough, written out a line at a time with narration explaining each step as it appears. Pause anywhere and ask why a line is there. Free to watch

Or read the worked solution below. Jump to the steps ↓

Written solution

P(C∣K)=1,P(C∣G)=12,P(G∣C)=16P(C\mid K)=1,\quad P(C\mid G)=\dfrac{1}{2},\quad P(G\mid C)=\dfrac{1}{6}

To find: P(K)\text{To find: } P(K)

K and G are complementary⇒P(K)+P(G)=1K \text{ and } G \text{ are complementary} \Rightarrow P(K)+P(G)=1

K:knows,G:guesses,C:correctK: \text{knows},\quad G: \text{guesses},\quad C: \text{correct}

P(K)=p,P(G)=1−pP(K)=p,\quad P(G)=1-p

P(C∣K)=1,P(C∣G)=12,P(W∣G)=12P(C\mid K)=1,\quad P(C\mid G)=\dfrac{1}{2},\quad P(W\mid G)=\dfrac{1}{2}

P(C)=P(C∣K) P(K)+P(C∣G) P(G)P(C)=P(C\mid K)\,P(K)+P(C\mid G)\,P(G)

Plan: 1. P(C)=P(C∣K)P(K)+P(C∣G)P(G)\text{Plan: } 1.\ P(C)=P(C\mid K)P(K)+P(C\mid G)P(G)

2. P(G∣C)=P(C∣G) P(G)P(C)2.\ P(G\mid C)=\dfrac{P(C\mid G)\,P(G)}{P(C)}

3. set P(G∣C)=16, find p3.\ \text{set } P(G\mid C)=\dfrac{1}{6},\ \text{find } p

P(C)=1⋅p+12(1−p)P(C)=1\cdot p+\dfrac{1}{2}(1-p)

P(C)=1+p2P(C)=\dfrac{1+p}{2}

P(G∣C)=P(C∣G) P(G)P(C)P(G\mid C)=\dfrac{P(C\mid G)\,P(G)}{P(C)}

P(G∣C)=12(1−p)1+p2=1−p1+pP(G\mid C)=\dfrac{\tfrac{1}{2}(1-p)}{\tfrac{1+p}{2}}=\dfrac{1-p}{1+p}

1−p1+p=16\dfrac{1-p}{1+p}=\dfrac{1}{6}

6(1−p)=1+p6(1-p)=1+p

6−6p=1+p6-6p=1+p

5=7p ⇒ p=575=7p\ \Rightarrow\ p=\dfrac{5}{7}

p=57 ⇒ P(G)=27p=\dfrac{5}{7}\ \Rightarrow\ P(G)=\dfrac{2}{7}

1−p1+p=2/712/7=16\dfrac{1-p}{1+p}=\dfrac{2/7}{12/7}=\dfrac{1}{6}

P(K)=57 ⇒ option (3)\boxed{P(K)=\dfrac{5}{7}\ \Rightarrow\ \text{option }(3)}

More from Probability (Class 12)

All 38 Probability (Class 12) questions →